Willard's age is eight less than three times Frank's age. The sum of their ages is 48. Find the age of each person.
Answer by Cromlix(4381)

Willard's age is eight less than three times Frank's age.
W = 3F - 8
The sum of their ages is 48.
F + W = 48
Replace W with 3F - 8 from
top equation.
F + 3F - 8 = 48
4F = 48 + 8
4F = 56
F = 14
Frank is 14 years old
Willard is 34 years old.
Hope this helps.
;-)